OpenStudy (calculusxy):

How to solve systems of linear and quadratic equations... -y^2 - 3x - 10y + 6 = 0 x + 3y - 2 = 0

OpenStudy (xapproachesinfinity):

substitution is you way out

OpenStudy (xapproachesinfinity):

have you tried?

OpenStudy (anonymous):

from second x=-3y+2 put in first \[-y^2-10y-3(-3y+2)+6=0\] \[-y^2-10y+9y-6+6=0\ find values of y and then corresponding x.

OpenStudy (calculusxy):

ok give me a moment please

OpenStudy (calculusxy):

so i got \(-y^2 + 6 + 9y - 10y + 6\) \(-y^2 - 1y + 12\)

OpenStudy (calculusxy):

then i would find the values of y by factoring it?

OpenStudy (xapproachesinfinity):

yeah quadratic equation you can use any method you are most familiar with

OpenStudy (anonymous):

\[-y^2+y=0,y(-y+1)=0,y=?\]

OpenStudy (calculusxy):

y = 4 y = -3

OpenStudy (calculusxy):

sorry i meant y = 3

OpenStudy (anonymous):

y=0,1

OpenStudy (calculusxy):

but how?

OpenStudy (phi):

you distributed -3 * +2 and got +6

OpenStudy (calculusxy):

okay so would i get now \(-y^2 - y\)

OpenStudy (phi):

equal to 0

OpenStudy (calculusxy):

ok

OpenStudy (calculusxy):

now i use the quadratic formula right?

OpenStudy (anonymous):

yes iwrote wrong it should be -y^2-y=0

OpenStudy (phi):

you could multiply both sides by -1 to get \[ y^2+y=0\] now factor out a y

OpenStudy (calculusxy):

\(y(y + 1)\)

OpenStudy (phi):

=0 you know either y is 0 or the factor (y+1) is 0

OpenStudy (anonymous):

take- y common ,then y=0,-1 find corresponding x from second eq.

OpenStudy (calculusxy):

okay so i would need to find what value for both of the y that will make it equal to 0

OpenStudy (phi):

are you asking how to solve y(y+1)=0 ?

OpenStudy (phi):

or, are you asking is there a different x value for each of the solutions for y ? (answer is yes, each y gives a different x)